ΔT = i Kbm Suppose we wanted to check the molecular weights of two sugars, glucose and sucrose, and we knew that sodium chloride the boiling point elevation for 1 m KCl solution. Suppose we have 10 g of each sugar. Then could we determine the molecular weights of the sugars via a boiling point elevation experiment?

Question 8 options:

Yes, the boiling point would be elevated twice as much for 1 mol of either sugar as for 1 mol KCl

No, we could not determine the molar mass because the sugars do not dissociate, so i=0. There could not be any boiling point elevation

The boiling point elevation would be the same no matter what salt we compared it with, so we could not determing anything

Yes, the boiling point would be elevated half as much for 1 mol of either sugar as for 1 mol KCl

No, we would need to know the ratio of the molar masses of the two sugars
Answer #1

Yes we can do it by the experiment.

i = vant Hoff factor = number of solute species present in solution = 1 for molecules which do not dissociate and = 2 for KCl because it dissociates to K+ and Cl-.

Kb = Ebullioscopic constant = constant for water.

m = molal concentration

Now, if the m = 1 then KCl gives a VI value twice that of glucose due to 'i' factor. So, if we divide the VI value of KCl solution by 2, then we'll get VI of 1 molal glucose solution. Now, again we'll dissolve 10 g of glucose in measured mass of water (say, M). So, calculated molality will be, moles of glucose in 1 Kg of water,

\frac{\frac{10\, g}{M_{glucose}\, g/mol}}{M\, g}\times \frac{1000\, g}{1\, Kg}=\frac{10000\, M}{M_{glucose}}\, molal

Mglucose is molar mass of glucose in g/mol.

Again, Since,

\Delta T\propto m

From the ratio of elevation of boiling point we'll get the actual molality of the glucose solution.

Then comparing these two we can calculate Mglucose. The same process applies for sucrose also.

So, the key point is, 1 m KCl solution has twice value of VI as compared to 1 m glucose or, sucrose. So, the correct option is the first one.